### v2

# REP: Resource-Efficient Prompting for Rehearsal-Free Continual Learning

Abstract
Recent rehearsal-free methods, guided by prompts, generally excel in vision-related continual learning (CL) scenarios with continuously drifting data. To be deployable on real-world devices, these methods must maintain high resource efficiency during training. In this paper, we introduce Resource-Efficient Prompting (REP), which targets improving the resource efficiency of prompt-based rehearsal-free methods. Our key focus is on avoiding catastrophic trade-offs with accuracy while trimming computational and memory costs during prompt learning. We achieve this by exploiting swift prompt selection that enhances input data using a carefully provisioned model, and by developing adaptive token merging (AToM) and layer dropping (ALD) algorithms for the prompt updating stage. AToM and ALD perform selective skipping across the data and model-layer dimensions without compromising task-specific features while learning new tasks. We validate REP’s superior resource efficiency over current state-of-the-art ViT- and CNN-based methods through extensive experiments on three image classification datasets.
1 Introduction
Continual learning (CL) trains neural network models on multiple sequential tasks, where each task may include data that diverges from previously encountered data. A crucial challenge for any CL algorithm is to effectively address catastrophic forgetting [22]. Severe forgetting occurs when a model rapidly loses previously learned knowledge while adapting to new tasks. The inability to adapt to new tasks not only impairs overall model accuracy but also affects user experiences of AI services in the real world [37].
Moreover, many of today’s AI services are designed for on-device scenarios [18, 10, 39] to securely learn tasks locally, without dependence on remote servers over the network. In on-device CL, learning methods must be optimized for computational (equivalently, energy) and memory efficiency requirements. Energy is scarce and can be intermittent in small devices in the wild [17, 14, 1], yet it is quickly drained by GPU operations during training [21]. Thus, achieving higher computational efficiency while training a new task in CL invariably enhances device durability. In contrast, memory efficiency often acts as a hard constraint. Device memory capacity is typically very limited and unified to operate all CPU and GPU programs. Therefore, a CL task that exhausts all available memory risks system crashes due to out-of-memory errors. As mobile and edge devices adhere to small form factors, it is difficult to dramatically augment battery and memory capacity to address these issues in practice.
In this paper, we present several new techniques for reducing the computational and memory costs of prompt-based rehearsal-free methods that leverage frozen, pre-trained vision transformer (ViT) models [35, 36, 34, 29, 15]. Prompts are a small set of parameters that progressively learn incoming tasks to combat forgetting. Updates to these compact prompts incur minimum data writes without significantly harming the lifespan of device storage, which typically sustains up to 10K writes per location. This makes using prompts a good fit for on-device CL. Moreover, thanks to recent advancements in small-memory ViT models, such as ViT-Ti [30] (5.8M parameters; similar size to ResNet-10), we can always spot powerful ViT models that match the resource efficiency of traditional CNN models under consideration across memory specifications of typical resource-constrained IoT and embedded devices, as shown in Table 1.
Figure 1: Overview of the proposed resource-efficient prompting (REP) algorithm for rehearsal-free CL. REP calculates query features from input samples using a lightweight model (e.g., ViT-Ti) to swiftly extract prompts from the prompt pool. These prompts are then inserted into a main backbone model (e.g., ViT-L) for training, which prioritizes model accuracy.
To enhance prompt-based CL for resource efficiency, we introduce REP (Resource-Efficient Prompting; Figure 1). REP is built upon our analysis of cost-accuracy trade-offs across the end-to-end learning process, with two key design insights. (1) The prompt selection stage, which constructs a prompt subset to augment input or intermediate data, is highly amenable to numerous promising options with fast approximations. (2) In contrast, the prompt update stage, which involves forward-backward passes over the backbone model, presents a range of optimizations with vastly different cost-accuracy trade-offs. With these insights, we develop three complementary techniques for both stages that trade minimal accuracy for high resource efficiency: (1) model downsizing for prompt selection and (2) adaptive token merging (AToM) and adaptive layer dropping (ALD) for prompt update. In pre-trained ViT models, task-specific knowledge is concentrated in shallow layers, while generalized feature information is spread across all layers [35] (See Figure 2 for our analysis). So, AToM and ALD carry out selective skipping across the data and model-layer dimensions in a “non-uniform” manner to preserve critical task-specific features in the shallow layers. REP provides resource efficiency that aligns with the capabilities of real-world commodity devices including NVIDIA Jetson TX2 and Nano. We evaluate REP across various memory budgets, backbone models, and datasets. Our experiments demonstrate that REP reduces wall-clock training time by 16–47% and lowers memory consumption by 5–33%.
2 Related Work
Memory Capacity ViT-Based Methods CNN-Based Methods
4-8 GB L2P (ViT-L), DualP (ViT-L) MEMO (RN26, RN34)
CODA-P (ViT-B) BudgetCL (RN101)
1-4 GB L2P (ViT-B), DualP (ViT-B) MEMO (RN10, RN18)
BudgetCL (RN34)
Up to 1 GB L2P (ViT-Ti), DualP (ViT-Ti) BudgetCL (RN10, RN18)
Table 1: ViT- and CNN-based methods with specific backbone models are mapped to the memory capacity they can fit into. DualP is DualPrompt, and CODA-P is CODA-Prompt.
Class-Incremental Learning. Our work primarily targets class-incremental learning (CIL) among various CL scenarios. In CIL, each incoming task introduces new classes within the same domain during the training phase, while task ID is not at hand during the inference phase [8]. In prior art using CNN backbone networks, rehearsal-based methods, which archive and replay representative samples from previous tasks [2, 4, 5, 28], have been a popular choice due to their superior performance [25]. Several recent methods are noteworthy along this line. For instance, BudgetCL [26] and CarM [19] suggest data-driven methods that outperform other strategies based on algorithmic optimizations, particularly when training is constrained by compute budgets. MEMO [40] effectively trades memory used to store old samples for saving task-specific model layers. We compare prompt-based methods with some of these methods and confirm better performance under resource constraints.
Continual Learning for the Edge. There has been a notable emphasis on efficient memory and energy usage in on-device learning, particularly in non-CL settings [9, 32, 33]. More recently, a handful of studies have explored extending CL capabilities to edge devices. Hayes and Kanan investigated online CL methods using CNN models tailored for embedded devices and provided valuable algorithmic insights. Similarly, Kwon et al. undertook a comparative analysis of rehearsal vs. regularization methods to uncover the cost-accuracy trade-offs, including storage, compute, and memory requirements. Ma et al. proposed an ML platform called Miro, which dynamically configures key design parameters of CNN-based CL methods to reduce energy costs within the device’s memory capacity. Unfortunately, none of these studies address the specific challenges associated with enabling vision transformers for on-device CL, which stands as a main goal of our research.
Prompting for Continual Learning. The use of prompts, which are explicit instructions or queries given to the model during training or inference, has shown promise in guiding continual learning for vision transformers [35, 36, 34, 29, 15, 13]. The general concept is to fine-tune prompts for new tasks while retaining knowledge acquired from previous tasks. L2P [36] and DualPrompt [35] propose prompt-tuning and prefix-tuning approaches, respectively, by selecting applicable prompts from a shared prompt pool. CODA-Prompt [29] introduces attention-conditioned prompts, which inherently facilitate higher prompting capacity. OVOR [13] uses only a single prompt to expedite the prompt selection stage. Our work suggests three complementary prompting techniques that can benefit all the above methods through full or even partial integration.
3 REP: Resource-Efficient Prompting
Figure 2: Mean attention distances for frozen blocks along (a) layer IDs and (b/c) attention heads. We run the first task of Split ImageNet-R. (a/b) L2P with ViT-L, and (c) DualPrompt with ViT-B.
Prompt-based rehearsal-free CL typically involves two key stages, prompt selection and prompt update. Figure 1 illustrates an overview of our method REP for making these stages more resource-efficient.
3.1 Prompt Selection
The prompt selection stage operates a neural network model $f_{\text{query}}$ to extract representations of input data from a prompt pool $P$. For a given input $x_{i}^{j}$ in task $T_{i}$, $f_{\text{query}}$ calculates a query feature $q(x_{i}^{j})\in\mathbb{R}^{D}$ and selects the prompt $p^{*}$ that maximizes the cosine similarity between the query and pre-trained prompt $p_{k}\in P$:
$p^{*}$ $=\underset{p_{k}\in P}{\mathrm{argmax}}\ \frac{\langle q(x_{i}^{j}),p_{k}\rangle}{\|q(x_{i}^{j})\|\|p_{k}\|}.$ (1)
Although $f_{\text{query}}$ is used only for model inference, its size, typically equivalent to the primary backbone model, can incur a substantial increase in computation time and memory usage. This is particularly true when it is implemented with a large transformer model. To mitigate these costs, we propose to adopt a more resource-efficient model $f_{\text{efficient}}$, which has a smaller depth and width than $f_{\text{query}}$.
Simply decreasing one dimension of the model can easily compromise the performance. Thus, we opt for a small pre-trained ViT model (ViT-Ti) as $f_{\text{efficient}}$, which we have found proficient in extracting essential representations without excessively cutting down on depth or width. When $f_{\text{efficient}}$ generates a query feature $q_{\text{efficient}}(x_{i}^{j})\in\mathbb{R}^{d}$ (where $d\leq D$), we apply a nonlinear random projection $\phi$ [23] to maintain the integrity of high-dimensional features in the selected prompt $p^{*}_{\text{efficient}}$, as follows:
$p_{\text{efficient}}^{*}$ $=\underset{p_{k}\in P}{\mathrm{argmax}}\ \frac{\langle\phi(q_{\text{efficient}}(x_{i}^{j})),p_{k}\rangle}{\|\phi(q_{\text{efficient}}(x_{i}^{j}))\|\|p_{k}\|}.$ (2)
Using ViT-Ti as $f_{\text{efficient}}$ enables a high level of feature overlap with $f_{\text{query}}$. For instance, on Split ImageNet-R, ViT-Ti achieves a prompt overlap of 76.3% on average with ViT-L, the largest $f_{\text{query}}$ in consideration, based on canonical-correlation analysis (CCA). This leads to nearly identical accuracy between $f_{\text{efficient}}$ and $f_{\text{query}}$ (refer to Table 2).
3.2 Prompt Update
The prompt update stage aims to refine learnable parameters for effective training and task adaptability by combining the softmax cross-entropy loss of the classifier $L_{\text{class}}$ and the cosine similarity loss of the prompt $L_{\text{prompt}}$. The combined loss function $L$ is expressed as:
$L$ $=L_{\text{class}}\left(f_{\text{update}}(x_{i}^{j}),y_{i}^{j}\right)+L_{\text{prompt}}\left(p^{*},q(x_{i}^{j})\right),$ (3)
where $f_{\text{update}}$ denotes the backbone model.
3.2.1 Unveiling System-Efficiency Opportunities
Remaining parameters that are not relevant to learnable parameters consist of a number of frozen transformer blocks. Then, are all frozen blocks useful for executing the loss function $L$, or can we skip some computations to improve resource efficiency?
To answer this question, we explore the feature representations of pre-trained transformer blocks with prompts by analyzing the mean attention distance. Mean attention distance reflects the degree to which an attention head focuses on a specific patch relative to others in the input. This metric is widely used to analyze the layer representation of ViT models during the pre-training process [27, 7]. In our study, we re-purpose this concept to examine the behavior of frozen layers during data drift. For a given head in an attention layer within frozen blocks, let $(x_{0},y_{0})$ denote the position of a query patch, and $(x_{i},y_{i})$ denote the positions of patches it attends to, with attention weights $a_{i}$. The distance $d_{i}$—typically defined as the Euclidean or pixel distance—is calculated as $d_{i}=(x_{i}-x_{0})^{2}+(y_{i}-y_{0})^{2}$. We then calculate the weighted mean distance as $\frac{\sum_{i}a_{i}\cdot d_{i}}{\sum_{i}a_{i}}$.
From a continual learning perspective, lower distances indicate a more localized attention pattern, thereby capturing “task-specific” features. Conversely, higher distances reflect a more global attention pattern, aligning with the extraction of more general, “task-agnostic” features. We measure mean attention distances along two dimensions: by layer ID and by attention head. Figure 2 shows these results when adapting to the first task of Split ImageNet-R.
In Figure 2(a), we present analysis results for the pre-trained layer ID using L2P with ViT-L. The lower layers exhibit highly dispersed attention distances, indicating that they simultaneously capture task-specific and task-agnostic information. However, as we examine deeper layers, these distances become increasingly concentrated at higher values, indicating that deeper layers specialize in capturing task-agnostic information. Taking a closer look at attention heads in a few selected layers in Figure 2(b), we observe that attention distances across the heads vary more widely in shallow layers than in deeper layers. Moreover, the lower distances observed in early heads of these shallow layers suggest that the learnable prompts integrated into the model effectively capture task-specific information. This behavior allows the model to retain task-relevant features while progressively specializing deeper layers for more generalized representations.
Similar behavior is observed when using DualPrompt with ViT-B in Fig. 2(c). Unlike L2P, DualPrompt integrates prompts directly into the input sequence before the attention layer of each transformer block, rather than only into the first block. This suggests that the feature representations of the pre-trained model, which balance local and global information, may not be greatly affected by the prompt method in use.
Next, we explain how we bring these insights into practice through two compute-skipping techniques.
3.2.2 Adaptive Token Merging (AToM)
Figure 3: The gradient norm for the prompt when applying conventional token merging (ToMe) vs AToM. L2P with ViT-L runs on Split ImageNet-R.
We first consider the data-efficient compute-skipping method via token merging. Conventional token merging (ToMe) [3] reduces the number of tokens by uniformly merging $n$ redundant tokens per layer, controlled by a fixed scheduler $r$. The scheduler function $r(l)\rightarrow n$ is applied to each layer $l$, and according to [3], it merges all tokens, including selected prompts. However, insights from Figure 2 and our additional analysis highlight two major problems.
First, there is a loss of task-specific information. The prompt tokens in CL carry essential task-specific information. However, ToMe indiscriminately combines these prompt tokens with others, thereby diminishing their intrinsic value. According to our empirical data in Figure 3, this approach can cause gradient explosions in the prompt tokens, even with gradient clipping, leading to learning instability. Second, there is a lack of layer-specific adaptability. Conventional token merging does not account for the disparity between shallow and deep layers, treating all layers uniformly. Therefore, there is a risk of excessive loss of valuable information from shallow layers, which are mainly responsible for adaptability to diverse sequential tasks.
Our adaptive token merging (AToM; Algorithm 1) addresses the loss of task-specific information by excluding prompt tokens during token merging, thus maintaining their specificity. To enhance task-specific adaptability, AToM uses a new scheduler $r^{\prime}(l)\rightarrow n^{\prime}$, which dynamically adjusts the number of tokens to merge based on layer depth, as follows:
$r^{\prime}(l)=\min(\delta\times(l-1),r_{\max}),$ (4)
where $l$ denotes the layer index, $\delta$ is the step change in the number of tokens to merge defined as $\frac{r_{\max}}{\text{L}-1}$ (with $L$ being the number of layers), and $r_{\max}$ is the maximum number of tokens to merge (by default, $2\times n$). With this $r^{\prime}(l)$, merging occurs more frequently in deeper layers than in shallow layers, preserving important task-related representations.
3.2.3 Adaptive Layer Dropping (ALD)
Figure 4: Comparing various layer dropping strategies using L2P with the ViT-L backbone on Split ImageNet-R.
Inspired by insights from Figure 2 and prior work on progressive layer dropping (PLD) [38], we propose adaptive layer dropping (ALD; Algorithm 2) with two key features: (1) the dropping schedule considers both temporal and spatial dimensions [^1] The spatial dimension refers to the different layers of the model. Each layer processes input features at varying levels of abstraction., and (2) it manages to drop layers non-uniformly to preserve critical task-specific features in shallower layers. On the contrary, PLD only considers the temporal aspect and does not differentiate between layers when dropping. This results in poorer performance compared to ALD, as shown in Table 3.
Algorithm 1 Adaptive Token Merging (AToM)
Input: Initial set of all tokens $T$; Set of prompt tokens $P$;
Number of model layers $L$;
Maximum number of tokens to merge $r_{\max}$
Initialize: $T^{\prime}_{\text{final}}\leftarrow T$;
1:  for $l\in\{1,2,\dots,\text{L}\}$ do
2:     $T_{\text{attn}}\leftarrow\text{MSA}(T^{\prime}_{\text{final}})$
3:     $T_{\text{eligible}}\leftarrow T_{\text{attn}}\setminus P$
4:     $\delta\leftarrow\frac{r_{\max}}{\text{L}-1}$
5:     $n^{\prime}\leftarrow\min(\delta\times(l-1),r_{\max})$
6:     $T^{\prime}_{\text{merged}}\leftarrow\text{Merge}(T_{\text{eligible}},n^{\prime})$
7:     $T_{\text{concat}}\leftarrow\text{Concat}(T^{\prime}_{\text{merged}},P)$
8:     $T^{\prime}_{\text{final}}\leftarrow\text{MLP}(T_{\text{concat}},l)$
9:  end for
10:  return $T^{\prime}_{\text{final}}$
To strengthen our claim, we conduct a simple experiment that drops 25% of layers from various positions. We compare their training wall-clock GPU time and final average accuracy in Figure 4 (See Section 4 for details on these metrics), using L2P with the ViT-L backbone on Split ImageNet-R. Specifically, Top-layer Drop and Bottom-layer Drop statically drop the first and last 25% of layers, respectively, while Random Drop randomly skips 25% of layers across the network. For comparison, we also include No Drop, which performs no layer dropping. Although all these naïve strategies incur accuracy loss, Bottom-layer Drop performs significantly worse than the others. Also, it appears undesirable to bypass only deep layers, even though these layers mainly capture general, task-agnostic features. While Random Drop most closely approaches No Drop’s accuracy, its uniform layer dropping still sacrifices some accuracy. This is why our method ALD carefully accounts for each layer’s spatial aspect in a non-uniform manner to minimize the impact on accuracy.
Algorithm 2 Adaptive Layer Dropping (ALD)
Input: Input tensor $X$; Keep ratio of layer $\theta_{t,l}$;
Number of layers $L$; Minimum ratio $\bar{\theta}$;
Decay rate $\gamma$; Spatial threshold $\tau$;
Adjustment factor $\alpha$
1:  for $l\in\{1,2,\dots,L\}$ do
2:     if $(n(l)-n^{\prime}(l))>\tau$ then
3:        $\alpha(l)\leftarrow\alpha$
4:     else
5:        $\alpha(l)\leftarrow 1$
6:     end if
7:     $\theta_{t,l}\leftarrow\alpha(l)\times\left((1-\bar{\theta})\exp(-\gamma\cdot t)+\bar{\theta}\right)$
8:     if $\text{Bernoulli}(\theta_{t,l})=1$ then
9:        $X_{\text{out}}\leftarrow\text{Exec}(l,X_{\text{out}})$
10:     else
11:        $X_{\text{out}}\leftarrow X_{\text{out}}$
12:     end if
13:  end for
14:  return $X_{\text{out}}$
ALD prioritizes retaining shallow layers that contain richer information essential for model performance, especially after token merging. Thus, instead of operating on its own schedule parameters, ALD leverages feedback from AToM, specifically the count of merged tokens at each layer, to guide layer dropping decisions. The layer-keeping probability $\theta_{t,l}$ is defined as:
$\theta_{t,l}=\left(\alpha(l)\times((1-\bar{\theta})\exp(-\gamma\cdot t)+\bar{\theta})\right),$ (5)
where $\theta_{t,l}$ is the probability of keeping layer $l$ at time step $t$, $\bar{\theta}$ is the minimum probability, $\gamma$ controls the rate of decay, and $\alpha(l)$ is the adjustment factor for layer $l$ defined as:
$\alpha(l)=\begin{cases}\alpha&\text{if }(n(l)-n^{\prime}(l))\geq\tau\\ 1&\text{if }(n(l)-n^{\prime}(l))<\tau.\end{cases}$ (6)
$\alpha(l)$ reflects the extent of token merging performed by AToM in layer $l$. $n(l)$ represents the original number of tokens, and $n^{\prime}(l)$ is the number of tokens remaining after merging. When the number of merged tokens surpasses the threshold $\tau$, ALD adjusts the layer-dropping probability based on $\alpha$. Since deeper layers tend to merge more tokens with AToM, ALD is more likely to exceed $\tau$ in deeper layers and drop more aggressively. These parameters should be tuned to balance between efficiency and the need to process the nuanced information contained in the merged tokens. We set $\alpha$ to 0.9, and $\tau$ to 16 for ViT-L, 12 for ViT-B, respectively.
One might contemplate exploiting stochastic depth [12], which gradually increases dropping probabilities for deeper layers without a temporal schedule. We compared ALD and stochastic depth and found that ALD generally offers 1.8–3.3% higher accuracy. This may be attributed to early training iterations after task insertion, which play a critical role in stabilizing training losses [21].
Resource Efficiency. REP triggers AToM and ALD for each task insertion to optimize both time and memory usage. Much like our prompt selection optimization, AToM consistently enhances time and memory efficiency. In contrast, ALD solely contributes to reducing time costs because it operates layer-keeping probability $\theta_{t,l}$ starting at 1.0, i.e., no layer dropping, which necessitates traversing full layers.
4 Experiments of REP
Figure 5: Energy-accuracy trade-offs of various ViT- and CNN-based methods over three different memory budgets: up to 1GB, 1-4GB, and 4-8. The memory breakdown of each method is on the first row. Experiments on Split CIFAR-100 and Split ImageNet-R are on the second and third row, respectively. ViT-based methods consistently outperform CNN-based methods.
We focus on the popular disjoint-CIL setup, where tasks consist of distinct sets of non-overlapping classes and samples from old classes are not given in future tasks [28, 4, 8].
Throughout this section, we will first describe the superiority of ViT-based prompting methods. However, it is important to note that the best-performing method is not fixed and can dynamically change depending on the device memory budget at hand. Therefore, to enhance prompt-based CL for better resource efficiency, we need adaptive techniques tailored to any backbone that function regardless of the specific learning method in use. REP is designed for this purpose, and we will evaluate this aspect in detail.
Data Generation. To organize task streams, we use three image classification datasets: CIFAR-100 (100 classes) [16], ImageNet-R (200 classes) [11], and PlantDisease (38 classes) [24]. Out of these datasets, ImageNet-R is known for exhibiting much higher intra-class variability among images and an uneven class-size distribution. We divide CIFAR-100 and ImageNet-R into 10 tasks to create Split CIFAR-100 (i.e., 10 classes per task) and Split ImageNet-R (i.e., 20 classes per task), respectively [36, 35, 29]. For PlantDisease, we drop 3 plant disease classes with very few images and organize the remaining 35 classes into 7 tasks to create Split PlantDisease (i.e., 5 classes per task).
Methods. We implement L2P [36], DualPrompt [35], and CODA-Prompt [29] as representative ViT-based prompting methods using PyTorch. L2P and DualPrompt select suitable prompts via inference, similar to [31], while CODA-Prompt employs a strategy for adaptively optimizing prompts without prompt selection, similar to [13]. These methods capitalize on ImageNet pre-trained models as backbones: ViT-L (307M), ViT-B (86M), and ViT-Ti (5.8M)—ViT-Ti is one of the smallest vision transformer models to our knowledge. We categorize the prompt-based methods into three groups based on their memory consumption, as detailed in Table 1, to reflect diverse device memory capacities.
To demonstrate the effectiveness of prompting methods, we also incorporate two state-of-the-art CNN-based methods, BudgetCL [6] and MEMO [40]. Both BudgetCL and MEMO improve model accuracy by leveraging spare memory to store and replay past samples, with MEMO additionally storing model checkpoints from history. These methods also utilize ImageNet pre-trained models as backbones: ResNet-10 (RN10; 5M), ResNet-18 (RN18; 11M), ResNet-26 (RN26; 14M), ResNet-34 (RN34; 22M), and ResNet-101 (RN101; 43M) models. For fair comparisons, we similarly group methods based on their memory usage. We refer readers to supplementary materials for detailed settings of all methods mentioned above.
Hardware and Metrics. Given that not all CL methods are currently feasible on small devices due to memory limitations, we use NVIDIA RTX 3090 Ti as a reference GPU and use the wall-clock time required to complete all tasks—Training wall-clock time correlates linearly with energy usage [21]. In supplementary materials, we evaluate our memory-efficient method directly on NVIDIA Jetson TX2, although it is applicable beyond this specific device. We also report the final average accuracy by averaging the accuracy of all classes after the last task training, and the forgetting metric [5, 29], which assesses the decline in task accuracy averaged over all tasks.
Figure 6: REP over the best methods across various memory budgets and datasets.
4.1 Preliminary Empirical Study
We first uncover the energy-accuracy trade-offs to highlight the impact of each baseline method on the accuracy gained relative to its training cost. Figure 5 visually represents the comparison results across device memory capacities and datasets. In each graph, the x-axis indicates wall-clock time, which corresponds to energy cost (lower is better), while the y-axis indicates final average accuracy (higher is better). Thus, a more cost-effective method appears closer to the upper-left corner of the graph. Overall, ViT-based methods outperform CNN-based methods by a large margin, achieving 26–36% higher accuracy with 45–90% less time and energy spent under the same memory budget. We elaborate on two key characteristics that explain these results:
$\bullet$ Performance scaling. ViT-based methods scale well, with larger backbone networks overall yielding higher accuracy. In contrast, CNN-based methods scale poorly with increased backbone size. Specifically, on Split ImageNet-R, a more challenging dataset, MEMO/RN34 is only 2.95% better than MEMO/RN18, despite being around twice larger. On Split CIFAR-100, increasing the backbone size yields only limited gains.
$\bullet$ Memory efficiency. CNN-based methods are often considered more memory-efficient because they use backbones a magnitude smaller than ViT models. Thus, it is commonly believed that ample memory can be allocated to replay samples or past model checkpoints [40], which help improve model performance. However, in CL, which involves typical training iterations, a substantial amount of memory must be allocated for feature maps [20], leaving little room for memory buffers. For instance, for a device memory range of 1–4GB in Table 1, the backbone of MEMO/RN18 consumes only 43MB of memory (6.5% of ViT-L), but feature maps occupy as much as 1.7GB (more than ViT-based methods). This causes MEMO/RN18 to be short on memory for the replay buffer, dramatically degrading model accuracy, as observed in  Figure 5.
4.2 REP on Best-Performing Methods
Next, in each graph in Figure 6, we demonstrate the efficacy of REP when applied to the best-performing CL method for each combination of memory budget and dataset observed in Figure 5. Accuracy is indicated by stars ($\bigstar$), while GPU wall-clock time is indicated by bars ($\blacksquare$). The memory usage in GB is noted on top of the bars. Please refer to the supplementary document for the full results of all datasets.
Resource Efficiency. The best-performing baselines optimized with REP demand significantly smaller amounts of system resources. Specifically, REP reduces memory consumption by 33% for L2P/ViT-L, 11% for DualPrompt/ViT-B, and 5% for DualPrompt/ViT-Ti. Despite the frozen backbone not requiring gradients for parameter updates, it still performs the backward pass to calculate classification and prompt losses for updating learnable parameters. Therefore, all intermediate data are stored in memory, similar to training a non-frozen model. This memory usage is reduced by applying AToM and ALD. In addition, REP exhibits a 16–47% reduction in GPU wall-clock time compared to competing methods, leading to great energy savings.
Accuracy. L2P, DualPrompt, and CODA-Prompt were initially evaluated using only ViT-B backbone, focusing primarily on accuracy in their original studies. In comparison, our evaluation extends these methods across diverse backbones while taking into account their energy-accuracy trade-offs in on-device CL across various memory constraints. When augmented with our techniques, these methods show significant improvements in resource efficiency with minimal accuracy loss across all memory budgets: 0.09–1.90% for Split PlantDisease, 0.03–1.07% for Split CIFAR-100, and 0.26–0.86% for Split ImageNet-R. Especially for the Split PlantDieases, REP with DualPrompt/ViT-B performed best under both 1–4GB and 4–8GB budgets.
4.3 Ablation and Additional Study
Ablated components Acc. ($\uparrow$) Fgt. ($\downarrow$) GPU Time (s) Mem. (GB)
REP (ours) 75.34$\pm$0.86 3.58$\pm$0.29 2542.43 4.45
(1) $f_{\text{efficient}}$ 74.79$\pm$0.78 4.04$\pm$0.80 3698.78 5.53
(2) AToM 74.91$\pm$0.77 3.47$\pm$0.84 4442.04 5.60
(3) ALD 74.53$\pm$0.78 3.63$\pm$0.67 4243.97 6.67
(4) $f_{\text{efficient}}$ + AToM 74.15$\pm$0.82 4.03$\pm$0.76 2861.35 4.46
(5) $f_{\text{efficient}}$ + ALD 74.52$\pm$0.43 3.43$\pm$0.54 3206.74 5.51
(6) AToM + ALD 74.57$\pm$0.37 2.56$\pm$0.52 3448.76 4.77
Table 2: Ablating REP’s components. They contribute to resource efficiency and accuracy.
We validate our proposed techniques and algorithm designs along with the chosen hyperparameters. The study here primarily uses the ViT-L backbone on Split ImageNet-R, where L2P is the best non-optimized method.
Acc. ($\uparrow$) Fgt. ($\downarrow$) GPU Time (s) Mem. (GB)
REP (ours) 75.34$\pm$0.86 3.58$\pm$0.29 2542.43 4.45
(1) w/ ToMe 70.19$\pm$0.74 2.63$\pm$0.92 2917.84 3.74
(2) w/ PLD 73.33$\pm$0.71 3.88$\pm$0.69 2741.90 4.46
Table 3: Accuracy of using traditional token merging (ToMe) or layer dropping (PLD) instead of our adaptive algorithms.
n (w/ $\theta$=0.5) Acc. ($\uparrow$) Fgt. ($\downarrow$) GPU Time (s) Mem. (GB)
2 76.16$\pm$0.74 3.54$\pm$0.68 2808.43 5.48
4 75.32$\pm$0.32 3.33$\pm$0.45 2714.04 5.16
8 75.34$\pm$0.86 3.58$\pm$0.29 2542.43 4.45
$\theta$ (w/ n=8) Acc. ($\uparrow$) Fgt. ($\downarrow$) GPU Time (s) Mem. (GB)
0.25 73.18$\pm$0.55 4.03$\pm$0.49 2362.07 4.45
0.50 75.34$\pm$0.86 3.58$\pm$0.29 2542.43 4.45
0.75 74.44$\pm$0.60 3.49$\pm$0.53 2861.35 4.45
Table 4: REP over different # of merged tokens ($n$) and % of keep ratio ($\theta$). We use $n$ = 8 and $\theta$ = 0.5 as default values.
Component Ablation. In  Table 2, we ablate REP ’s components to assess their contributions to resource efficiency. Each component plays a key role in reducing memory and compute time, with ALD affecting compute time only, as expected. Notably, AToM substantially optimizes both resource usages.
Algorithm Validation. Token merging (ToMe) [3] and progressive layer dropping (PLD) [38] are specifically designed to accelerate traditional transformer-based model training. To validate the importance of incorporating our adaptive techniques instead in CL, we evaluate how REP performs in case AToM and ALD are replaced with ToMe and PLD, respectively. The results are presented in Table 3. Although applying ToMe or PLD improves system efficiency over L2P, it results in an accuracy loss of 5.15% or 2.01%, respectively, compared to using our techniques. This indicates that ToMe and PLD are less desirable for achieving our goal, i.e., enhancing system efficiency without compromising accuracy.
AToM and ALD Intensity. Table 4 shows the effects of varying intensities of AToM and ALD, focusing on the number of merged tokens ($n$) in AToM and the keep ratio ($\theta$) in ALD. AToM appears to maintain stable accuracy even as more tokens are merged. In contrast, for ALD, a lower keep ratio improves system efficiency by reducing iteration time and memory usage, but it can markedly impair model accuracy if the ratio is too low. Overall, when used with caution, AToM and ALD can effectively balance system efficiency and accuracy.
Longer Task Sequence. We examine how REP performs on a longer task sequence of 20 tasks. A longer task sequence represents a more challenging scenario, where performance drop as the number of tasks increases can become more pronounced. We observe that REP consistently preserves model accuracy (from 72.77$\pm$0.19 in L2P to 72.58$\pm$0.59 in REP) while improving resource efficiency by reducing GPU wall-clock time and memory usage by 47.3% and 33%, respectively.
Method Acc. ($\uparrow$) Fgt. ($\downarrow$) GPU Time (s) Mem. (GB)
REP (ours) 75.34$\pm$0.16 3.58$\pm$0.29 2542.43 4.45
L2P 61.57$\pm$0.66 9.73$\pm$0.47 1609.88 14.20
DualPrompt 68.13$\pm$0.49 4.68$\pm$0.20 1507.42 12.30
CODA-Prompt 75.45$\pm$0.56 1.64$\pm$0.10 1853.54 18.83
Table 5: Results on Split ImageNet-R, with ViT-B as the backbone model (following the original setups).
Competing Methods with Original Setups. For L2P, DualPrompt, and CODA-Prompt, we used the hyperparameters of the original papers, except for the number of iterations per task and batch size. In particular, smaller batch sizes must be adopted to fit into on-device memory constraints. Nonetheless, we present the results of three competing methods on Split ImageNet-R using their original hyperparameters in Table 5. Despite using the ViT-B backbone, they consume significantly more memory than REP due to their large batch sizes.
5 Discussion
Since our study focuses on scenarios wherein a sequence of tasks is presented within learning environments, our method does not determine the curriculum of learning. In general, actively selecting such curricula could enhance the final average accuracy, motivating future work to extend our method for curriculum-based continual learning. Moreover, although using replay buffer may not be feasible in real-world scenarios due to data privacy concerns, on-device CL is already viewed as a privacy-preserving approach, making rehearsal-based REP a promising direction.

### v3

@@
 Abstract
-Recent rehearsal-free methods, guided by prompts, generally excel in vision-related continual learning (CL) scenarios with continuously drifting data.
-To be deployable on real-world devices, these methods must maintain high resource efficiency during training.
-In this paper, we introduce Resource-Efficient Prompting (REP), which targets improving the resource efficiency of prompt-based rehearsal-free methods.
-Our key focus is on avoiding catastrophic trade-offs with accuracy while trimming computational and memory costs during prompt learning.
-We achieve this by exploiting swift prompt selection that enhances input data using a carefully provisioned model, and by developing adaptive token merging (AToM) and layer dropping (ALD) algorithms for the prompt updating stage.
-AToM and ALD perform selective skipping across the data and model-layer dimensions without compromising task-specific features while learning new tasks.
-We validate REP’s superior resource efficiency over current state-of-the-art ViT- and CNN-based methods through extensive experiments on three image classification datasets.
+Recent rehearsal-free methods, guided by prompts, excel in vision-related continual learning (CL) with drifting data but lack resource efficiency, making real-world deployment challenging.
+In this paper, we introduce Resource-Efficient Prompting (REP), which improves the computational and memory efficiency of prompt-based rehearsal-free methods while minimizing accuracy trade-offs.
+Our approach employs swift prompt selection to refine input data using a carefully provisioned model and introduces adaptive token merging (AToM) and layer dropping (ALD) for efficient prompt updates.
+AToM and ALD selectively skip data and model layers while preserving task-specific features during new task learning.
+Extensive experiments on multiple image classification datasets demonstrate REP’s superior resource efficiency over state-of-the-art ViT- and CNN-based methods.
 1 Introduction
@@
 As mobile and edge devices adhere to small form factors, it is difficult to dramatically augment battery and memory capacity to address these issues in practice.
-In this paper, we present several new techniques for reducing the computational and memory costs of prompt-based rehearsal-free methods that leverage frozen, pre-trained vision transformer (ViT) models [35, 36, 34, 29, 15].
+In this paper, we present several new techniques to reduce the computational and memory costs of prompt-based rehearsal-free methods that leverage frozen, pre-trained vision transformer (ViT) models [38, 39, 37, 31, 16].
 Prompts are a small set of parameters that progressively learn incoming tasks to combat forgetting.
@@
 (2) In contrast, the prompt update stage, which involves forward-backward passes over the backbone model, presents a range of optimizations with vastly different cost-accuracy trade-offs.
-With these insights, we develop three complementary techniques for both stages that trade minimal accuracy for high resource efficiency: (1) model downsizing for prompt selection and (2) adaptive token merging (AToM) and adaptive layer dropping (ALD) for prompt update.
+With these insights, we develop three complementary techniques for both stages that trade minimal accuracy drop for high resource efficiency: (1) model downsizing for prompt selection and (2) adaptive token merging (AToM) and adaptive layer dropping (ALD) for prompt update.
 In pre-trained ViT models, task-specific knowledge is concentrated in shallow layers, while generalized feature information is spread across all layers [38] (See Figure 2 for our analysis).
@@
 We evaluate REP across various memory budgets, backbone models, and datasets.
-Our experiments demonstrate that REP reduces wall-clock training time by 16–47% and lowers memory consumption by 5–33%.
+Our experiments demonstrate that REP reduces training iteration time by up to 52% and lowers memory consumption by up to 32%.
 2 Related Work
+Table 1: ViT- and CNN-based methods with specific backbone models are mapped to the memory capacity they can fit into. DualP is DualPrompt, and CODA-P is CODA-Prompt.
 Memory Capacity ViT-Based Methods CNN-Based Methods
@@
 Up to 1 GB L2P (ViT-Ti), DualP (ViT-Ti) BudgetCL (RN10, RN18)
-Table 1: ViT- and CNN-based methods with specific backbone models are mapped to the memory capacity they can fit into. DualP is DualPrompt, and CODA-P is CODA-Prompt.
-Class-Incremental Learning.
-Our work primarily targets class-incremental learning (CIL) among various CL scenarios.
+Class-Incremental Learning: Our work primarily targets class-incremental learning (CIL) among various CL scenarios.
 In CIL, each incoming task introduces new classes within the same domain during the training phase, while task ID is not at hand during the inference phase [8].
@@
 We compare prompt-based methods with some of these methods and confirm better performance under resource constraints.
-Continual Learning for the Edge.
-There has been a notable emphasis on efficient memory and energy usage in on-device learning, particularly in non-CL settings [9, 32, 33].
+Continual Learning for the Edge: There has been a notable emphasis on efficient memory and energy usage in on-device learning, particularly in non-CL settings [9, 35, 36].
 More recently, a handful of studies have explored extending CL capabilities to edge devices.
@@
 undertook a comparative analysis of rehearsal vs.
-regularization methods to uncover the cost-accuracy trade-offs, including storage, compute, and memory requirements.
+regularization methods to uncover cost-accuracy trade-offs, including storage, compute, and memory requirements.
 Ma et al.
-proposed an ML platform called Miro, which dynamically configures key design parameters of CNN-based CL methods to reduce energy costs within the device’s memory capacity.
-Unfortunately, none of these studies address the specific challenges associated with enabling vision transformers for on-device CL, which stands as a main goal of our research.
-Prompting for Continual Learning.
-The use of prompts, which are explicit instructions or queries given to the model during training or inference, has shown promise in guiding continual learning for vision transformers [35, 36, 34, 29, 15, 13].
+proposed Miro, an ML platform that dynamically configures key design parameters of CNN-based CL methods to reduce energy costs within the device’s memory capacity.
+However, none of these studies address the challenges associated with enabling vision transformers for on-device CL, which stands as a main goal of our research.
+Prompting for Continual Learning: The use of prompts, which are explicit instructions or queries given to the model during training or inference, has shown promise in guiding continual learning for vision transformers [38, 39, 37, 31, 16, 14].
 The general concept is to fine-tune prompts for new tasks while retaining knowledge acquired from previous tasks.
@@
 In our study, we re-purpose this concept to examine the behavior of frozen layers during data drift.
-For a given head in an attention layer within frozen blocks, let $(x_{0},y_{0})$ denote the position of a query patch, and $(x_{i},y_{i})$ denote the positions of patches it attends to, with attention weights $a_{i}$.
+For a given head in an attention layer within frozen blocks, let $(x_{0},y_{0})$ denote the position of a query patch, and $(x_{i},y_{i})$ the positions of patches that the attention head attends to, with corresponding attention weights $a_{i}$.
 The distance $d_{i}$—typically defined as the Euclidean or pixel distance—is calculated as $d_{i}=(x_{i}-x_{0})^{2}+(y_{i}-y_{0})^{2}$.
@@
 3.2.2 Adaptive Token Merging (AToM)
-Figure 3: The gradient norm for the prompt when applying conventional token merging (ToMe) vs AToM. L2P with ViT-L runs on Split ImageNet-R.
 We first consider the data-efficient compute-skipping method via token merging.
@@
 Therefore, there is a risk of excessive loss of valuable information from shallow layers, which are mainly responsible for adaptability to diverse sequential tasks.
+Figure 3: The norm of gradient with respect to the prompt while training Split ImageNet-R when AToM and ToMe (Conventional Token Merging) is applied. ViT-L with L2P is used as a backbone.
 Our adaptive token merging (AToM; Algorithm 1) addresses the loss of task-specific information by excluding prompt tokens during token merging, thus maintaining their specificity.
@@
 Figure 4: Comparing various layer dropping strategies using L2P with the ViT-L backbone on Split ImageNet-R.
-Inspired by insights from Figure 2 and prior work on progressive layer dropping (PLD) [38], we propose adaptive layer dropping (ALD; Algorithm 2) with two key features: (1) the dropping schedule considers both temporal and spatial dimensions [^1] The spatial dimension refers to the different layers of the model.
-Each layer processes input features at varying levels of abstraction., and (2) it manages to drop layers non-uniformly to preserve critical task-specific features in shallower layers.
+Inspired by insights from Figure 2 and prior work on progressive layer dropping (PLD) [41], we propose adaptive layer dropping (ALD; Algorithm 2) with two key features: (1) the dropping schedule considers both temporal and spatial dimensions [^1] The spatial dimension refers to model layers, each processing input features at different levels of abstraction., and (2) it manages to drop layers non-uniformly to preserve critical task-specific features in shallower layers.
 On the contrary, PLD only considers the temporal aspect and does not differentiate between layers when dropping.
 This results in poorer performance compared to ALD, as shown in Table 4.
+To strengthen our claim, we conduct a simple experiment that drops 25% of layers from various positions.
+We compare their training wall-clock GPU time and final average accuracy in Figure 4, using L2P with the ViT-L backbone on Split ImageNet-R.
+Specifically, Top-layer Drop and Bottom-layer Drop statically drop the first and last 25% of layers, respectively, while Random Drop randomly skips 25% of layers across the network.
+For comparison, we also include No Drop, which performs no layer dropping.
+Although all these naïve strategies incur accuracy loss, Top-layer Drop performs significantly worse than the others.
+Also, it appears undesirable to bypass only deep layers, even though these layers mainly capture general, task-agnostic features.
+While Random Drop most closely approaches No Drop’s accuracy, its uniform layer dropping still sacrifices some accuracy.
+This is why our method ALD carefully accounts for each layer’s spatial aspect in a non-uniform manner to minimize the impact on accuracy.
+ALD prioritizes retaining shallow layers that contain richer information essential for model performance, especially after token merging.
+Thus, instead of operating on its own schedule parameters, ALD leverages feedback from AToM, specifically the count of merged tokens at each layer, to guide layer dropping decisions.
+The layer-keeping probability $\theta_{t,l}$ is defined as:
+$\theta_{t,l}=\left(\alpha(l)\times((1-\bar{\theta})\exp(-\gamma\cdot t)+\bar{\theta})\right),$ (5)
+where $\theta_{t,l}$ is the probability of keeping layer $l$ at time step $t$, $\bar{\theta}$ is the minimum probability, $\gamma$ controls the rate of decay, and $\alpha(l)$ is the adjustment factor for layer $l$ defined as:
+$\alpha(l)=\begin{cases}\alpha&\text{if }(n(l)-n^{\prime}(l))\geq\tau\\ 1&\text{if }(n(l)-n^{\prime}(l))<\tau.\end{cases}$ (6)
+$\alpha(l)$ quantifies the degree of token merging performed by AToM in layer $l$.
+$n(l)$ represents the original number of tokens, and $n^{\prime}(l)$ is the number of tokens remaining after merging.
+When the number of merged tokens surpasses the threshold $\tau$, ALD adjusts the layer-dropping probability based on $\alpha$.
+Since deeper layers tend to merge more tokens with AToM, ALD is more likely to exceed $\tau$ in deeper layers and drop more aggressively.
+These parameters should be tuned to balance between efficiency and preserving the nuanced information contained in the merged tokens.
+We set $\alpha$ to 0.9, and $\tau$ to 16 for ViT-L, 12 for ViT-B, respectively.
+One might contemplate exploiting stochastic depth [13], which gradually increases dropping probabilities for deeper layers without a temporal schedule.
+We compared ALD and stochastic depth and found that ALD generally offers 1.8–3.3% higher accuracy.
+This may be attributed to early training iterations after task insertion, which help stabilize training losses [23].
+Resource Efficiency: REP triggers AToM and ALD for each task insertion to optimize both time and memory usage.
+Much like our prompt selection optimization, AToM consistently enhances time and memory efficiency.
+In contrast, ALD solely contributes to reducing time costs because it operates layer-keeping probability $\theta_{t,l}$ starting at 1.0, i.e., no layer dropping, which necessitates traversing full layers.
+Figure 5: Energy-accuracy trade-offs of various ViT- and CNN-based methods over three different memory budgets: up to 1GB, 1-4GB, and 4-8.
+The memory breakdown of each method is on the first row.
+Experiments on Split CIFAR-100 and Split ImageNet-R are on the second and third row, respectively.
+ViT-based methods consistently outperform CNN-based methods.
+4 Experiments
+We focus on the popular disjoint-CIL setup, where tasks consist of distinct sets of non-overlapping classes and samples from old classes are not given in future tasks [30, 4, 8].
+Throughout this section, we will first describe the superiority of ViT-based prompting methods.
+However, it is important to note that the best-performing method is not fixed and can dynamically change depending on the device memory budget at hand.
+Therefore, to enhance prompt-based CL for better resource efficiency, we need adaptive techniques tailored to any backbone that function regardless of the specific learning method in use.
+REP is designed for this purpose, and we will evaluate this aspect in detail.
+Data Generation: To organize task streams, we use three image classification datasets: CIFAR-100 (100 classes) [18], ImageNet-R (200 classes) [12], and PlantDisease (38 classes) [26].
+Out of these datasets, ImageNet-R is known for exhibiting much higher intra-class variability among images and an uneven class-size distribution.
+We divide CIFAR-100 and ImageNet-R into 10 tasks to create Split CIFAR-100 (i.e., 10 classes per task) and Split ImageNet-R (i.e., 20 classes per task), respectively [39, 38, 31].
+For PlantDisease, we drop 3 plant disease classes with very few images and organize the remaining 35 classes into 7 tasks to create Split PlantDisease (i.e., 5 classes per task).
+Methods: We employ L2P [39], DualPrompt [38], and CODA-Prompt [31] as representative ViT-based prompting methods.
+L2P and DualPrompt select suitable prompts via inference, similar to [34], while CODA-Prompt employs a strategy for adaptively optimizing prompts without prompt selection, similar to [14].
+These methods capitalize on ImageNet pre-trained models as backbones: ViT-L (307M), ViT-B (86M), and ViT-Ti (5.8M)—ViT-Ti is one of the smallest vision transformer models to our knowledge.
+We categorize the prompt-based methods into these three backbone models.
+To demonstrate the effectiveness of prompting methods, we also incorporate two state-of-the-art CNN-based methods, BudgetCL [6] and MEMO [43].
+Both BudgetCL and MEMO improve model accuracy by leveraging spare memory to store and replay past samples, with MEMO additionally storing model checkpoints from history.
+These methods also use ImageNet pre-trained models as backbones: ResNet-10 (RN10; 5M), ResNet-18 (RN18; 11M), ResNet-26 (RN26; 14M), ResNet-34 (RN34; 22M), and ResNet-101 (RN101; 43M) models.
+For fair comparisons, we similarly group methods based on their memory usage.
+The analysis result is presented in  Section 4.1.
+See  Appendix B for detailed settings of all methods mentioned above.
+Hardware and Metrics: We use NVIDIA RTX 3090 as a reference GPU to cover a wide range of memory sizes while maintaining consistent computational power and measure the GPU iteration time required for computation—GPU iteration time can be correlated linearly with energy usage [23].
+In Appendix H, we evaluate our memory-efficient method directly on NVIDIA Jetson TX2, although it is applicable beyond this specific device.
+We also report the final average accuracy by averaging the accuracy of all classes after the last task training [5, 31], which assesses the decline in task accuracy averaged over all tasks.
+4.1 Preliminary Empirical Study
+Table 2: Accuracy and computational cost of all competing methods on Split CIFAR-100, Split ImageNet-R, and Split PlantDisease datasets.
+We report Final Average Accuracy (Acc.) both without and with REP.
+Iteration time (ms) and memory usage (GB) are provided in absolute terms, along with their percentage values relative to REP.
+($\uparrow$) higher is better, ($\downarrow$) lower is better.
+Model Method Split CIFAR-100 Split ImageNet-R Split PlantDisease Iter. time (ms) Mem. (GB)
+Acc. ($\uparrow$) w/ REP Acc. ($\uparrow$) w/ REP Acc. ($\uparrow$) w/ REP Iter.time ($\downarrow$) w/ REP Mem. ($\downarrow$) w/ REP
+ViT-L L2P 88.2$\pm$0.3 86.4$\pm$0.3 75.6$\pm$1.0 75.3$\pm$0.9 75.9$\pm$3.1 81.1$\pm$3.9 446.9 235.4 (-47.3%) 6.5 4.5 (-31.5%)
+DualPrompt 86.3$\pm$0.3 84.5$\pm$0.6 71.2$\pm$0.6 70.6$\pm$0.9 75.1$\pm$1.1 78.2$\pm$2.9 423.7 202.5 (-52.2%) 5.9 4.3 (-27.8%)
+CODA-Prompt 85.4$\pm$0.4 83.7$\pm$0.8 74.4$\pm$0.6 72.4$\pm$0.9 75.0$\pm$2.0 74.0$\pm$2.7 568.0 431.1 (-24.1%) 13.2 11.0 (-16.8%)
+UpperBound 94.1$\pm$0.1 - 85.2$\pm$0.2 - 99.6$\pm$0.3 - 426.7 - 9.8 -
+ViT-B L2P 84.5$\pm$0.8 83.4$\pm$1.8 59.4$\pm$0.8 59.7$\pm$1.2 64.4$\pm$2.6 64.0$\pm$2.4 142.5 100.1 (-29.8%) 2.3 2.0 (-11.8%)
+DualPrompt 85.1$\pm$0.2 84.9$\pm$0.9 68.1$\pm$0.1 67.0$\pm$1.0 78.0$\pm$1.5 77.0$\pm$1.8 133.4 99.7 (-25.2%) 2.1 1.8 (-11.2%)
+CODA-Prompt 83.3$\pm$0.9 81.3$\pm$1.9 59.0$\pm$0.3 58.0$\pm$1.7 71.4$\pm$0.8 70.2$\pm$2.7 164.1 144.8 (-11.8%) 6.9 5.7 (-17.8%)
+UpperBound 92.0$\pm$0.2 - 81.4$\pm$0.1 - 99.6$\pm$0.4 - 143.2 - 3.2 -
+ViT-Ti L2P 60.3$\pm$0.2 58.9$\pm$1.3 41.2$\pm$0.7 40.3$\pm$1.0 56.1$\pm$2.9 55.0$\pm$2.8 33.8 29.0 (-14.2%) 0.5 0.5 (-2.1%)
+DualPrompt 62.9$\pm$0.5 62.5$\pm$0.9 43.6$\pm$0.8 42.7$\pm$1.0 64.0$\pm$1.0 61.9$\pm$1.8 33.9 30.0 (-11.6%) 0.4 0.4 (-4.8%)
+CODA-Prompt 67.0$\pm$0.4 65.9$\pm$1.8 49.3$\pm$0.8 47.6$\pm$2.0 66.8$\pm$0.8 65.0$\pm$2.6 36.1 29.8 (-17.4%) 2.8 2.8 (-0.7%)
+UpperBound 83.0$\pm$0.1 - 77.4$\pm$0.2 - 98.0$\pm$0.9 - 42.3 - 0.6 -
+We first uncover the energy-accuracy trade-offs to highlight the impact of each baseline method on the accuracy gained relative to its training cost.
+Figure 5 visually represents the comparison results across device memory capacities and datasets.
+In each graph, the x-axis indicates wall-clock time, which corresponds to energy cost (lower is better), while the y-axis indicates final average accuracy (higher is better).
+Thus, a more cost-effective method appears closer to the upper-left corner of the graph.
+Overall, ViT-based methods outperform CNN-based methods by a large margin, achieving 26–36% higher accuracy with 45–90% less time and energy spent under the same memory budget.
+We elaborate on two key characteristics that explain these results:
+Scaling Across Varying Backbone Networks.
+ViT-based methods scale well, with larger backbone networks overall yielding higher accuracy.
+In contrast, CNN-based methods scale poorly with increased backbone size.
+Specifically, on Split ImageNet-R, a more challenging dataset, MEMO/RN34 is only 2.95% better than MEMO/RN18, despite being around twice larger.
+On Split CIFAR-100, increasing the backbone size yields only limited gains.
+Memory Efficiency.
+CNN-based methods are often considered more memory-efficient because they use backbones a magnitude smaller than ViT models.
+Thus, it is commonly believed that ample memory can be allocated to replay samples or past model checkpoints [43], which help improve model performance.
+However, in CL, which involves typical training iterations, a substantial amount of memory must be allocated for feature maps [22], leaving little room for memory buffers.
+For instance, for a device memory range of 1–4GB in Table 1, the backbone of MEMO/RN18 consumes only 43MB of memory (6.5% of ViT-L), but feature maps occupy as much as 1.7GB (more than ViT-based methods).
+This causes MEMO/RN18 to be short on memory for the replay buffer, dramatically degrading model accuracy, as observed in Figure 5.
+4.2 Main Results
+Table 2 presents the efficacy of REP when integrated into prompt-based CL methods across various ViT backbones (ViT-L, ViT-B, ViT-Ti) and datasets.
+Resource Efficiency:  Integrating REP into existing prompt-based CL methods significantly reduces computational overhead and memory usage across all model configurations: we observe memory savings of up to 32% for ViT-L, 18% for ViT-B, and 5% for ViT-Ti.
+Although the backbone network parameters remain frozen during training, the model must still perform a backward pass to compute losses for both classification and prompt learning.
+As a result, a substantial amount of memory is allocated to store intermediate activations.
+By selectively merging tokens (AToM) and adaptively dropping layers (ALD), REP mitigates these memory demands without compromising performance-critical information flow.
+The advantages of REP extend to training speed, which is closely linked to energy consumption.
+For ViT-L, the largest architecture considered, REP reduces iteration time by up to 52% compared to baseline methods, translating directly into lower energy consumption.
+In contrast, with the more compact ViT-Ti backbone, resource savings persist but are less pronounced due to the model’s already low computational cost.
+These results highlight the versatility of REP, demonstrating its scalability from compact to large models while ensuring substantial resource efficiency gains across different regimes.
+Accuracy:  While L2P, DualPrompt, and CODA-Prompt were originally designed and evaluated on a ViT-B backbone, our study systematically extends these methods across a diverse range of ViT architectures, considering their energy-accuracy trade-offs in on-device CL under various memory constraints.
+This broader perspective reveals notable gains in resource efficiency with REP, archived with only marginal accuracy drops, e.g., 0.03–1.36% on Split CIFAR-100 and 0.26–1.62% on Split ImageNet-R.
+In some cases, accuracy even improves.
+For instance, on ViT-L with Split PlantDisease, L2P’s accuracy increases from 75.92% to 81.08% when REP is applied.
+These findings suggest that AToM and ALD effectively eliminate redundant computations while preserving essential features for CL.
+Table 3: Ablation study of REP components, demonstrating their contributions to both resource efficiency and accuracy.
+Ablated components Acc. ($\uparrow$) Fgt. ($\downarrow$) GPU Time (s) Mem. (GB)
+REP (ours) 75.6$\pm$1.0 3.6$\pm$0.3 2542.4 4.5
+(1) $f_{\text{efficient}}$ 74.8$\pm$0.8 4.0$\pm$0.8 3698.8 5.5
+(2) AToM 74.9$\pm$0.8 3.5$\pm$0.8 4442.0 5.6
+(3) ALD 74.5$\pm$0.8 3.6$\pm$0.7 4244.0 6.7
+(4) $f_{\text{efficient}}$ + AToM 74.2$\pm$0.8 4.0$\pm$0.8 2861.4 4.5
+(5) $f_{\text{efficient}}$ + ALD 74.5$\pm$0.4 3.4$\pm$0.5 3206.7 5.5
+(6) AToM + ALD 74.6$\pm$0.4 2.6$\pm$0.5 3448.8 4.8
+4.3 Ablation and Additional Study
+We validate our proposed techniques and algorithm designs along with the chosen hyperparameters.
+The study here primarily uses the ViT-L backbone on Split ImageNet-R, where L2P is the best non-optimized method.
+Table 4: Accuracy of using traditional token merging (ToMe) or layer dropping (PLD) instead of our adaptive algorithms.
+Acc. ($\uparrow$) Fgt. ($\downarrow$) GPU Time (s) Mem. (GB)
+REP (ours) 75.6$\pm$1.0 3.6$\pm$0.3 2542.4 4.5
+(1) w/ ToMe 70.2$\pm$0.7 2.6$\pm$0.9 2917.8 3.7
+(2) w/ PLD 73.3$\pm$0.7 3.9$\pm$0.7 2741.9 4.5
+Table 5: REP over different # of merged tokens ($n$) and % of keep ratio ($\theta$). We use $n$ = 8 and $\theta$ = 0.5 as default values.
+$n$ (w/ $\theta$=0.5) Acc. ($\uparrow$) Fgt. ($\downarrow$) GPU Time (s) Mem. (GB)
+2 76.2$\pm$0.7 3.5$\pm$0.7 2808.4 5.5
+4 75.3$\pm$0.3 3.3$\pm$0.5 2714.0 5.2
+8 75.6$\pm$0.9 3.6$\pm$0.3 2542.4 4.5
+$\theta$ (w/ $n$=8) Acc. ($\uparrow$) Fgt. ($\downarrow$) GPU Time (s) Mem. (GB)
+0.25 73.2$\pm$0.6 4.0$\pm$0.5 2362.1 4.5
+0.50 75.6$\pm$1.0 3.6$\pm$0.3 2542.4 4.5
+0.75 74.4$\pm$0.6 3.5$\pm$0.5 2861.4 4.5
+Component Ablation: In  Table 3, we ablate REP ’s components to assess their contributions to resource efficiency.
+Each component plays a key role in reducing memory and compute time, with ALD affecting compute time only, as expected.
+Notably, AToM substantially optimizes both resource usages.
+Algorithm Validation: Token merging (ToMe) [3] and progressive layer dropping (PLD) [41] are specifically designed to accelerate traditional transformer-based model training.
+To validate the importance of incorporating our adaptive techniques instead in CL, we evaluate how REP performs in case AToM and ALD are replaced with ToMe and PLD, respectively.
+The results are presented in Table 4.
+Although applying ToMe or PLD improves system efficiency over L2P, it results in an accuracy loss of 5.15% or 2.01%, respectively, compared to using our techniques.
+This indicates that ToMe and PLD are less desirable for achieving our goal, i.e., enhancing system efficiency without compromising accuracy.
+AToM and ALD Intensity: Table 5 shows the effects of varying intensities of AToM and ALD, focusing on the number of merged tokens ($n$) in AToM and the keep ratio ($\theta$) in ALD.
+AToM appears to maintain stable accuracy even as more tokens are merged.
+In contrast, for ALD, a lower keep ratio improves system efficiency by reducing iteration time and memory usage, but it can markedly impair model accuracy if the ratio is too low.
+Overall, when used with caution, AToM and ALD can effectively balance system efficiency and accuracy.
+5 Conclusion
+In this work, we introduced Resource-Efficient Prompting (REP), a framework designed to enhance the computational and memory efficiency of prompt-based rehearsal-free continual learning methods.
+By incorporating adaptive token merging and adaptive layer dropping, REP selectively streamlines the learning process while preserving task-specific features, significantly reducing resource consumption without compromising accuracy.
+Experiments on multiple image classification datasets demonstrate that REP achieves substantial efficiency gains over state-of-the-art ViT- and CNN-based methods, making it a practical solution for on-device continual learning.
+Since our study focuses on scenarios wherein a sequence of tasks is presented within learning environments, our method does not determine the curriculum of learning.
+Future work could extend our our method for curriculum-based continual learning.
+Moreover, although using replay buffer may not be feasible in real-world scenarios due to data privacy concerns, on-device CL is already viewed as a privacy-preserving approach, making rehearsal-based REP a promising direction.
+Appendix A Algorithm Details
 Algorithm 1 Adaptive Token Merging (AToM)
@@
 10:  return $T^{\prime}_{\text{final}}$
-To strengthen our claim, we conduct a simple experiment that drops 25% of layers from various positions.
-We compare their training wall-clock GPU time and final average accuracy in Figure 4 (See Section 4 for details on these metrics), using L2P with the ViT-L backbone on Split ImageNet-R.
-Specifically, Top-layer Drop and Bottom-layer Drop statically drop the first and last 25% of layers, respectively, while Random Drop randomly skips 25% of layers across the network.
-For comparison, we also include No Drop, which performs no layer dropping.
-Although all these naïve strategies incur accuracy loss, Bottom-layer Drop performs significantly worse than the others.
-Also, it appears undesirable to bypass only deep layers, even though these layers mainly capture general, task-agnostic features.
-While Random Drop most closely approaches No Drop’s accuracy, its uniform layer dropping still sacrifices some accuracy.
-This is why our method ALD carefully accounts for each layer’s spatial aspect in a non-uniform manner to minimize the impact on accuracy.
 Algorithm 2 Adaptive Layer Dropping (ALD)
@@
 14:  return $X_{\text{out}}$
-ALD prioritizes retaining shallow layers that contain richer information essential for model performance, especially after token merging.
-Thus, instead of operating on its own schedule parameters, ALD leverages feedback from AToM, specifically the count of merged tokens at each layer, to guide layer dropping decisions.
-The layer-keeping probability $\theta_{t,l}$ is defined as:
-$\theta_{t,l}=\left(\alpha(l)\times((1-\bar{\theta})\exp(-\gamma\cdot t)+\bar{\theta})\right),$ (5)
-where $\theta_{t,l}$ is the probability of keeping layer $l$ at time step $t$, $\bar{\theta}$ is the minimum probability, $\gamma$ controls the rate of decay, and $\alpha(l)$ is the adjustment factor for layer $l$ defined as:
-$\alpha(l)=\begin{cases}\alpha&\text{if }(n(l)-n^{\prime}(l))\geq\tau\\ 1&\text{if }(n(l)-n^{\prime}(l))<\tau.\end{cases}$ (6)
-$\alpha(l)$ reflects the extent of token merging performed by AToM in layer $l$.
-$n(l)$ represents the original number of tokens, and $n^{\prime}(l)$ is the number of tokens remaining after merging.
-When the number of merged tokens surpasses the threshold $\tau$, ALD adjusts the layer-dropping probability based on $\alpha$.
-Since deeper layers tend to merge more tokens with AToM, ALD is more likely to exceed $\tau$ in deeper layers and drop more aggressively.
-These parameters should be tuned to balance between efficiency and the need to process the nuanced information contained in the merged tokens.
-We set $\alpha$ to 0.9, and $\tau$ to 16 for ViT-L, 12 for ViT-B, respectively.
-One might contemplate exploiting stochastic depth [12], which gradually increases dropping probabilities for deeper layers without a temporal schedule.
-We compared ALD and stochastic depth and found that ALD generally offers 1.8–3.3% higher accuracy.
-This may be attributed to early training iterations after task insertion, which play a critical role in stabilizing training losses [21].
-Resource Efficiency.
-REP triggers AToM and ALD for each task insertion to optimize both time and memory usage.
-Much like our prompt selection optimization, AToM consistently enhances time and memory efficiency.
-In contrast, ALD solely contributes to reducing time costs because it operates layer-keeping probability $\theta_{t,l}$ starting at 1.0, i.e., no layer dropping, which necessitates traversing full layers.
-4 Experiments of REP
-Figure 5: Energy-accuracy trade-offs of various ViT- and CNN-based methods over three different memory budgets: up to 1GB, 1-4GB, and 4-8.
-The memory breakdown of each method is on the first row.
-Experiments on Split CIFAR-100 and Split ImageNet-R are on the second and third row, respectively.
-ViT-based methods consistently outperform CNN-based methods.
-We focus on the popular disjoint-CIL setup, where tasks consist of distinct sets of non-overlapping classes and samples from old classes are not given in future tasks [28, 4, 8].
-Throughout this section, we will first describe the superiority of ViT-based prompting methods.
-However, it is important to note that the best-performing method is not fixed and can dynamically change depending on the device memory budget at hand.
-Therefore, to enhance prompt-based CL for better resource efficiency, we need adaptive techniques tailored to any backbone that function regardless of the specific learning method in use.
-REP is designed for this purpose, and we will evaluate this aspect in detail.
-Data Generation.
-To organize task streams, we use three image classification datasets: CIFAR-100 (100 classes) [16], ImageNet-R (200 classes) [11], and PlantDisease (38 classes) [24].
-Out of these datasets, ImageNet-R is known for exhibiting much higher intra-class variability among images and an uneven class-size distribution.
-We divide CIFAR-100 and ImageNet-R into 10 tasks to create Split CIFAR-100 (i.e., 10 classes per task) and Split ImageNet-R (i.e., 20 classes per task), respectively [36, 35, 29].
-For PlantDisease, we drop 3 plant disease classes with very few images and organize the remaining 35 classes into 7 tasks to create Split PlantDisease (i.e., 5 classes per task).
-Methods.
-We implement L2P [36], DualPrompt [35], and CODA-Prompt [29] as representative ViT-based prompting methods using PyTorch.
-L2P and DualPrompt select suitable prompts via inference, similar to [31], while CODA-Prompt employs a strategy for adaptively optimizing prompts without prompt selection, similar to [13].
-These methods capitalize on ImageNet pre-trained models as backbones: ViT-L (307M), ViT-B (86M), and ViT-Ti (5.8M)—ViT-Ti is one of the smallest vision transformer models to our knowledge.
-We categorize the prompt-based methods into three groups based on their memory consumption, as detailed in Table 1, to reflect diverse device memory capacities.
-To demonstrate the effectiveness of prompting methods, we also incorporate two state-of-the-art CNN-based methods, BudgetCL [6] and MEMO [40].
-Both BudgetCL and MEMO improve model accuracy by leveraging spare memory to store and replay past samples, with MEMO additionally storing model checkpoints from history.
-These methods also utilize ImageNet pre-trained models as backbones: ResNet-10 (RN10; 5M), ResNet-18 (RN18; 11M), ResNet-26 (RN26; 14M), ResNet-34 (RN34; 22M), and ResNet-101 (RN101; 43M) models.
-For fair comparisons, we similarly group methods based on their memory usage.
-We refer readers to supplementary materials for detailed settings of all methods mentioned above.
-Hardware and Metrics.
-Given that not all CL methods are currently feasible on small devices due to memory limitations, we use NVIDIA RTX 3090 Ti as a reference GPU and use the wall-clock time required to complete all tasks—Training wall-clock time correlates linearly with energy usage [21].
-In supplementary materials, we evaluate our memory-efficient method directly on NVIDIA Jetson TX2, although it is applicable beyond this specific device.
-We also report the final average accuracy by averaging the accuracy of all classes after the last task training, and the forgetting metric [5, 29], which assesses the decline in task accuracy averaged over all tasks.
-Figure 6: REP over the best methods across various memory budgets and datasets.
-4.1 Preliminary Empirical Study
-We first uncover the energy-accuracy trade-offs to highlight the impact of each baseline method on the accuracy gained relative to its training cost.
-Figure 5 visually represents the comparison results across device memory capacities and datasets.
-In each graph, the x-axis indicates wall-clock time, which corresponds to energy cost (lower is better), while the y-axis indicates final average accuracy (higher is better).
-Thus, a more cost-effective method appears closer to the upper-left corner of the graph.
-Overall, ViT-based methods outperform CNN-based methods by a large margin, achieving 26–36% higher accuracy with 45–90% less time and energy spent under the same memory budget.
-We elaborate on two key characteristics that explain these results:
-$\bullet$ Performance scaling.
-ViT-based methods scale well, with larger backbone networks overall yielding higher accuracy.
-In contrast, CNN-based methods scale poorly with increased backbone size.
-Specifically, on Split ImageNet-R, a more challenging dataset, MEMO/RN34 is only 2.95% better than MEMO/RN18, despite being around twice larger.
-On Split CIFAR-100, increasing the backbone size yields only limited gains.
-$\bullet$ Memory efficiency.
-CNN-based methods are often considered more memory-efficient because they use backbones a magnitude smaller than ViT models.
-Thus, it is commonly believed that ample memory can be allocated to replay samples or past model checkpoints [40], which help improve model performance.
-However, in CL, which involves typical training iterations, a substantial amount of memory must be allocated for feature maps [20], leaving little room for memory buffers.
-For instance, for a device memory range of 1–4GB in Table 1, the backbone of MEMO/RN18 consumes only 43MB of memory (6.5% of ViT-L), but feature maps occupy as much as 1.7GB (more than ViT-based methods).
-This causes MEMO/RN18 to be short on memory for the replay buffer, dramatically degrading model accuracy, as observed in  Figure 5.
-4.2 REP on Best-Performing Methods
-Next, in each graph in Figure 6, we demonstrate the efficacy of REP when applied to the best-performing CL method for each combination of memory budget and dataset observed in Figure 5.
-Accuracy is indicated by stars ($\bigstar$), while GPU wall-clock time is indicated by bars ($\blacksquare$).
-The memory usage in GB is noted on top of the bars.
-Please refer to the supplementary document for the full results of all datasets.
-Resource Efficiency.
-The best-performing baselines optimized with REP demand significantly smaller amounts of system resources.
-Specifically, REP reduces memory consumption by 33% for L2P/ViT-L, 11% for DualPrompt/ViT-B, and 5% for DualPrompt/ViT-Ti.
-Despite the frozen backbone not requiring gradients for parameter updates, it still performs the backward pass to calculate classification and prompt losses for updating learnable parameters.
-Therefore, all intermediate data are stored in memory, similar to training a non-frozen model.
-This memory usage is reduced by applying AToM and ALD.
-In addition, REP exhibits a 16–47% reduction in GPU wall-clock time compared to competing methods, leading to great energy savings.
-Accuracy.
-L2P, DualPrompt, and CODA-Prompt were initially evaluated using only ViT-B backbone, focusing primarily on accuracy in their original studies.
-In comparison, our evaluation extends these methods across diverse backbones while taking into account their energy-accuracy trade-offs in on-device CL across various memory constraints.
-When augmented with our techniques, these methods show significant improvements in resource efficiency with minimal accuracy loss across all memory budgets: 0.09–1.90% for Split PlantDisease, 0.03–1.07% for Split CIFAR-100, and 0.26–0.86% for Split ImageNet-R.
-Especially for the Split PlantDieases, REP with DualPrompt/ViT-B performed best under both 1–4GB and 4–8GB budgets.
-4.3 Ablation and Additional Study
+Appendix B Implementation Details
+Unless otherwise stated, we use the term REP-L to denote REP applied to L2P with a ViT-L backbone, which stands out as the best-performing baseline method for Split CIFAR-100 and Split ImageNet-R within the 4–8GB memory range.
+Similarly, the notations L, B, and Ti are used to generally denote ViT-L, ViT-B, and ViT-Ti backbones, respectively, in the prompt update stage of any prompt-based method.
+B.1 Prompt-Based Methods for REP
+We present details on the three major prompt-based methods to which REP is integrated for resource efficiency.
+Learning to Prompt (L2P) positions prompts at the first layer of the transformer architecture [39].
+These prompts are learnable parameters that dynamically evolve with the training process.
+The mechanism begins with a prompt pool $P=\{p_{1},p_{2},...,p_{m}\}\subset\mathbb{R}^{L_{p}\times D}$.
+For a given input $x_{i}^{j}$ in task $T_{i}$, L2P computes a query feature $q(x_{i}^{j})\in\mathbb{R}^{D}$ to select the corresponding prompt.
+The prompt $p^{*}$ is selected based on maximizing the cosine similarity with respect to the query:
+$p^{*}=\underset{p_{k}\in P}{\mathrm{argmax}}\sum_{c=1}^{L_{p}}\frac{\big{\langle}q(x_{i}^{j}),[p_{k}]_{(c,:)}^{\top}\big{\rangle}}{\|q(x_{i}^{j})\|\|[p_{k}]_{(c,:)}^{\top}\|},$ (7)
+where $[p_{k}]_{(c,:)}$ is the $c$-th row of $p_{k}$.
+The selected prompt $p^{*}$ is concatenated with the input embedding $z(x_{i}^{j})$ to form the prompt-augmented input $z^{\prime}(x_{i}^{j})=[p^{*};z(x_{i}^{j})]$.
+The training objective of L2P balances classification loss $L_{\text{class}}$ and prompt-adjustment loss $L_{\text{prompt}}$:
+$L=L_{\text{class}}(z^{\prime}(x_{i}^{j}),y_{i}^{j})+L_{\text{prompt}}(p^{*},q(x_{i}^{j})).$ (8)
+DualPrompt leverages prompts at multiple layers of the transformer architecture [38].
+It introduces a general prompt $g\in\mathbb{R}^{L_{g}\times D}$ and a set of task-specific prompts $E=\{e_{1},e_{2},...,e_{T}\}\subset\mathbb{R}^{L_{e}\times D}$.
+These prompts are incorporated at specified layers in the transformer model.
+For a given input $x_{i}^{j}$ in task $T_{i}$, the model’s transformer layers $f$ are modified by attaching $g$ and $e_{i}$ to the layers, resulting in a prompted architecture $f_{g,e_{i}}$.
+The feature transformation $h_{i}^{j}$ for the input sample $x_{i}^{j}$ is then obtained as:
+$h_{i}^{j}=f_{g,e_{i}}(x_{i}^{j}).$ (9)
+Similar to L2P, DualPrompt optimizes $L_{\text{class}}$ and $L_{\text{prompt}}$:
+$L=L_{\text{class}}(h_{i}^{j},y_{i}^{j})+L_{\text{prompt}}(g,e_{i}).$ (10)
+CODA-Prompt decomposes learnable prompts into components and uses an attention mechanism from a pre-trained ViT model to select relevant prompts [31].
+Instead of a single prompt, CODA-Prompt learns a set of prompt components $P=\{P_{1},P_{2},...,P_{M}\}$.
+The final prompt $p$ is a weighted sum:
+$p=\sum_{m}\alpha_{m}P_{m},$ (11)
+where weights $\alpha$ are determined based on the query $q(x)$ and keys $K\in\mathbb{R}^{D\times M}$:
+$\alpha=\gamma(q(x),K).$ (12)
+When the task changes, the current components are frozen, and new one is added, promoting orthogonality:
+$L_{\text{ortho}}(B)=\|BB^{\top}-I\|^{2},$ (13)
+where $B$ represents $P$, $K$, or $A$ ($A$ represents the attention vector). The full optimization target is:
+$\min_{P^{n},K^{n},A^{n},\phi^{n}}$ $\,L\left(f_{\phi}(f_{\theta,P,K,A}(x)),y\right)+\lambda\left(L_{\text{ortho}}(P)+L_{\text{ortho}}(K)+L_{\text{ortho}}(A)\right)$ (14)
+where $P^{n}$, $K^{n}$, and $A^{n}$ are new components, and $\lambda$ balances the orthogonality loss [31].
+B.2 Hyperparameters and Configurations
+For all prompt-based CL methods, we employ the ADAM optimizer [17] configured with hyperparameters $\beta_{1}=0.99$ and $\beta_{2}=0.999$.
+In contrast, BudgetCL and MEMO adopt the stochastic gradient descent (SGD) optimizer with a momentum coefficient of 0.9 and a weight decay of 0.002, both validated as effective by similar replay buffer-based methods [39].
+In REP, we set the prompt pool size to 10 and the prompt length to 5, following the original implementations of L2P [^2] L2P PyTorch Implementation.
+https://github.com/JH-LEE-KR/l2p-pytorch.
+and DualPrompt.
+[^3] DualPrompt PyTorch Implementation.
+https://github.com/JH-LEE-KR/dualprompt-pytorch.
+The learning rate for the prompt-based methods is set to 0.001875.
+For CODA-Prompt, we implement a cosine decay learning rate strategy, as outlined in its original paper [31].
+For BudgetCL and MEMO, the learning rate is set to 0.1 and decays by 0.1 at 80 and 150 epochs following the original paper [43].
+We consistently use the mini-batch size of 16 for all methods to maintain a uniform computational load for each training iteration.
+For fair comparisons, each method performs 1875, 1080, and 2583 iterations per task insertion for Split CIFAR-100 [18], Split ImageNet-R [12], and Split PlantDisease [26], respectively.
+This setup ensures that among prompt-based methods, iteration time itself correlates linearly with energy usage, as training wall-clock time similarly reflects energy consumption as discussed in the main paper.
+BudgetCL and MEMO select 8 samples from the new task and another 8 samples from old tasks in the replay buffer to organize 16 samples in each mini-batch.
+We adopt the setup in [43] and train 200 epochs for all datasets used.
+Thus, BudgetCL and MEMO perform more iterations per epoch for a bigger replay buffer, translating into longer GPU time and more energy usage.
+Appendix C Competing Methods with Original Setups
+Table 6: Results on Split CIFAR-100, with ViT-B as the backbone model (following the original setups).
+Method Acc. ($\uparrow$) Fgt. ($\downarrow$) GPU Time (s) Mem. (GB)
+REP-L (ours) 86.89$\pm$0.57 4.65$\pm$0.14 4420.99 4.45
+L2P-B 84.36$\pm$0.68 7.35$\pm$0.38 2802.42 14.20
+DualPrompt-B 85.27$\pm$0.75 5.16$\pm$0.09 2624.06 12.30
+CODA-Prompt-B 86.09$\pm$0.73 1.67$\pm$0.26 3226.58 18.83
+Table 7: Results on Split ImageNet-R, with ViT-B as the backbone model (following the original setups).
+Method Acc. ($\uparrow$) Fgt. ($\downarrow$) GPU Time (s) Mem. (GB)
+REP-L (ours) 75.34$\pm$0.16 3.58$\pm$0.29 2542.43 4.45
+L2P-B 59.66$\pm$0.81 9.73$\pm$0.47 1609.88 14.20
+DualPrompt-B 67.45$\pm$0.88 4.68$\pm$0.20 1507.42 12.30
+CODA-Prompt-B 75.52$\pm$0.74 1.64$\pm$0.10 1853.54 18.83
+For the evaluation of L2P, DualPrompt, and CODA-Prompt in the main paper, we adhere to the hyperparameters specified in the original studies, except for the number of iterations per task and batch size.
+We adopt a smaller batch size to account for on-device memory limitations, which prevent the use of the original batch sizes.
+Nonetheless, we present the results of these three methods using their original hyperparameters in Table 6 and Table 7, respectively.
+Although these methods are based on the ViT-B backbone, they consume significantly more memory than REP-L (with the ViT-L backbone), primarily due to their originally large batch sizes.
+Appendix D Additional Datasets
+Table 8: Results on Split CUB-200, with ViT-L as the backbone model.
+Method Acc. ($\uparrow$) Fgt. ($\downarrow$) GPU Time (s) Mem. (GB)
+REP-L (ours) 74.70$\pm$1.04 6.67$\pm$0.34 529.67 4.45
+L2P-L 74.70$\pm$1.02 6.68$\pm$1.35 1005.57 6.67
+DualPrompt-L 72.35$\pm$0.98 7.82$\pm$0.31 953.42 5.93
+CODA-Prompt-L 79.46$\pm$0.84 5.84$\pm$0.99 1257.82 13.95
+We also expand our experiments on REP by incorporating an additional dataset: Split CUB-200 [33], which is designed for fine-grained image classification.
+Split CUB-200 contains 5,994 bird images categorized into 200 classes, which are split into 5 CL tasks, each having 40 classes.
+We present the experimental results for all prompt-based methods using ViT-L as the backbone in Table 8.
+REP-L outperforms competing methods in resource efficiency by bringing significant reductions in both computation time and memory usage.
+Importantly, REP-L is on par with its baseline model L2P in accuracy.
+Note that while adopting CODA-Prompt to implement REP can deliver better accuracy, its memory usage far exceeds all memory ranges under consideration.
+Appendix E Extended Ablation and Additional Study
+As the paper mainly discusses the ablation study based on Split ImageNet-R, we here present an extended ablation study of REP using Split CIFAR-100.
+The results are shown in Table 9.
+Similar to the findings from Split ImageNet-R, the ablation of any component within REP for the Split CIFAR-100 dataset impacts resource efficiency.
+All components contribute to the reduction of computation time and memory usage, with AToM particularly standing out in both aspects of resource efficiency.
+Furthermore, we perform algorithm validation on Split CIFAR-100, for which we present the results in Table 10.
+Consistent with the Split ImageNet-R results, using static token merging (ToMe) [3] or progressive layer dropping (PLD) [41] instead of adaptive token merging (AToM) or adaptive layer dropping (ALD) negatively affects model accuracy on the Split CIFAR-100 dataset.
+This experiment further corroborates the efficacy of our proposed algorithms in optimizing resource usage without compromising model accuracy in CL scenarios.
+Table 9: Extended component ablation for Split CIFAR-100. Ablating any component of REP results in lower resource efficiency by increasing training time and memory consumption.
 Ablated components Acc. ($\uparrow$) Fgt. ($\downarrow$) GPU Time (s) Mem. (GB)
-REP (ours) 75.34$\pm$0.86 3.58$\pm$0.29 2542.43 4.45
-(1) $f_{\text{efficient}}$ 74.79$\pm$0.78 4.04$\pm$0.80 3698.78 5.53
-(2) AToM 74.91$\pm$0.77 3.47$\pm$0.84 4442.04 5.60
-(3) ALD 74.53$\pm$0.78 3.63$\pm$0.67 4243.97 6.67
-(4) $f_{\text{efficient}}$ + AToM 74.15$\pm$0.82 4.03$\pm$0.76 2861.35 4.46
-(5) $f_{\text{efficient}}$ + ALD 74.52$\pm$0.43 3.43$\pm$0.54 3206.74 5.51
-(6) AToM + ALD 74.57$\pm$0.37 2.56$\pm$0.52 3448.76 4.77
-Table 2: Ablating REP’s components. They contribute to resource efficiency and accuracy.
-We validate our proposed techniques and algorithm designs along with the chosen hyperparameters.
-The study here primarily uses the ViT-L backbone on Split ImageNet-R, where L2P is the best non-optimized method.
+REP-L (ours) 86.89$\pm$0.57 4.65$\pm$0.14 4420.99 4.45
+(1) $f_{\text{efficient}}$ 86.38$\pm$0.43 5.82$\pm$0.81 6096.14 5.53
+(2) AToM 86.43$\pm$0.23 4.94$\pm$0.77 7724.21 5.60
+(3) ALD 86.79$\pm$0.73 5.01$\pm$0.39 7379.79 6.67
+(4) $f_{\text{efficient}}$ + AToM 85.97$\pm$0.54 4.99$\pm$0.73 4975.57 4.46
+(5) $f_{\text{efficient}}$ + ALD 86.07$\pm$0.18 4.70$\pm$0.66 5576.16 5.51
+(6) AToM + ALD 86.08$\pm$0.48 3.97$\pm$0.42 5997.01 4.77
+Table 10: Extended algorithm validation for Split CIFAR-100. Changing our adaptive algorithms to conventional static token merging (ToMe) or progressive layer dropping (PLD) harms model accuracy.
 Acc. ($\uparrow$) Fgt. ($\downarrow$) GPU Time (s) Mem. (GB)
-REP (ours) 75.34$\pm$0.86 3.58$\pm$0.29 2542.43 4.45
-(1) w/ ToMe 70.19$\pm$0.74 2.63$\pm$0.92 2917.84 3.74
-(2) w/ PLD 73.33$\pm$0.71 3.88$\pm$0.69 2741.90 4.46
-Table 3: Accuracy of using traditional token merging (ToMe) or layer dropping (PLD) instead of our adaptive algorithms.
-n (w/ $\theta$=0.5) Acc. ($\uparrow$) Fgt. ($\downarrow$) GPU Time (s) Mem. (GB)
-2 76.16$\pm$0.74 3.54$\pm$0.68 2808.43 5.48
-4 75.32$\pm$0.32 3.33$\pm$0.45 2714.04 5.16
-8 75.34$\pm$0.86 3.58$\pm$0.29 2542.43 4.45
-$\theta$ (w/ n=8) Acc. ($\uparrow$) Fgt. ($\downarrow$) GPU Time (s) Mem. (GB)
-0.25 73.18$\pm$0.55 4.03$\pm$0.49 2362.07 4.45
-0.50 75.34$\pm$0.86 3.58$\pm$0.29 2542.43 4.45
-0.75 74.44$\pm$0.60 3.49$\pm$0.53 2861.35 4.45
-Table 4: REP over different # of merged tokens ($n$) and % of keep ratio ($\theta$). We use $n$ = 8 and $\theta$ = 0.5 as default values.
-Component Ablation.
-In  Table 2, we ablate REP ’s components to assess their contributions to resource efficiency.
-Each component plays a key role in reducing memory and compute time, with ALD affecting compute time only, as expected.
-Notably, AToM substantially optimizes both resource usages.
-Algorithm Validation.
-Token merging (ToMe) [3] and progressive layer dropping (PLD) [38] are specifically designed to accelerate traditional transformer-based model training.
-To validate the importance of incorporating our adaptive techniques instead in CL, we evaluate how REP performs in case AToM and ALD are replaced with ToMe and PLD, respectively.
-The results are presented in Table 3.
-Although applying ToMe or PLD improves system efficiency over L2P, it results in an accuracy loss of 5.15% or 2.01%, respectively, compared to using our techniques.
-This indicates that ToMe and PLD are less desirable for achieving our goal, i.e., enhancing system efficiency without compromising accuracy.
-AToM and ALD Intensity.
-Table 4 shows the effects of varying intensities of AToM and ALD, focusing on the number of merged tokens ($n$) in AToM and the keep ratio ($\theta$) in ALD.
-AToM appears to maintain stable accuracy even as more tokens are merged.
-In contrast, for ALD, a lower keep ratio improves system efficiency by reducing iteration time and memory usage, but it can markedly impair model accuracy if the ratio is too low.
-Overall, when used with caution, AToM and ALD can effectively balance system efficiency and accuracy.
-Longer Task Sequence.
-We examine how REP performs on a longer task sequence of 20 tasks.
-A longer task sequence represents a more challenging scenario, where performance drop as the number of tasks increases can become more pronounced.
-We observe that REP consistently preserves model accuracy (from 72.77$\pm$0.19 in L2P to 72.58$\pm$0.59 in REP) while improving resource efficiency by reducing GPU wall-clock time and memory usage by 47.3% and 33%, respectively.
+REP-L (ours) 86.89$\pm$0.57 4.65$\pm$0.14 4420.99 4.45
+(1) w/ ToMe 83.55$\pm$0.14 3.73$\pm$0.61 5073.80 3.74
+(2) w/ PLD 84.40$\pm$0.35 4.88$\pm$0.83 4767.86 4.46
+Appendix F Extended Preliminary Empirical Study
+Figure 6: Extended analysis of the energy-accuracy trade-off on Split PlantDisease dataset.
+While CNN-based methods may excel on certain memory ranges, ViT-based methods consistently demonstrate better resource efficiency.
+We present an extended analysis of the energy-accuracy trade-off using the Split PlantDisease dataset [26] in Figure 6.
+Interestingly, CNN-based methods can achieve higher accuracy on this dataset, albeit with higher computational costs.
+Since the PlantDisease dataset consists of localized leaf images with various disease patterns, it appears to favor CNN architectures like ResNets, which are excellent at capturing local features.
+On the contrary, ViT-based methods, which excel at capturing global information due to their pre-trained weights, may not perform very well on tasks requiring extensive localized feature extraction.
+Nonetheless, ViT-based methods demonstrate significantly greater resource efficiency.
+Thus, for this specific benchmark, the choice between CNN-based methods and ViT-base methods simply depends on the priority: higher resource efficiency vs.
+higher accuracy.
+Appendix G REP under Varying Compute Budgets
+Table 11: Impact of REP under various computation budgets (i.e., number of iterations per task).
+Method 313 625 938 1250 1563 1875
+REP-L (ours) 84.06 85.47 86.60 86.66 86.57 86.89
+L2P-L 84.36 86.28 87.34 87.76 87.83 88.24
+We compare REP-L with L2P across various compute budgets defined by the number of iterations per task.
+The comparison results using Split CIFAR-100 are shown in Table 11.
+Evidently, the accuracy of REP-L is comparable to that of L2P across six compute budgets, with the performance gap getting increasingly marginal at lower compute budgets.
+We observe that the overall accuracy for both methods does not drop significantly with reduced compute budgets.
+This is attributed to the fact that Split CIFAR-100 is considered a less complex CL benchmark.
+Appendix H REP on Edge Devices
+Miro [23] introduces a dynamic approach for fine-tuning key design parameters of rehearsal-based CL methods to achieve high model accuracy while simultaneously reducing energy costs, i.e., high cost-effectiveness.
+Miro’s methodology centers on identifying the optimal memory size within the device’s capacity to accommodate both old and new samples for training.
+We compare REP with Miro directly on a reference edge device NVIDIA Jetson TX2444NVIDIA.
+Jetson TX2 Module.
+https://developer.nvidia.com/embedded/
+jetson-tx2.
+.
+This device is equipped with a 256-core NVIDIA Pascal GPU, 8GB of RAM, and a 32GB eMMC 5.1 drive.
+The RAM is a single unified memory shared by the CPU and GPU.
+We report energy usage as joules (J) obtained by multiplying power by time.
+Power usage is measured by reading built-in sensor values provided by Jetson devices for the GPU, RAM, CPU, and I/O connection [^5] Convenient Power Measurements on the Jetson TX2/Tegra X2 Board.
+https://embeddeddl.wordpress.com/2018/04/25/convenient-power-measurements-on-the-jetson-tx2-tegra-x2-board/..
+H.1 Power Usage Breakdown
+Figure 7: Power breakdown for training a ViT-B model on NVIDIA Jetson TX2.
+When implementing CL on edge devices, the majority of power consumption that influences energy usage during the training of a new task comes from GPU operations.
+Figure 7 shows power consumption on NVIDIA Jetson TX2 across individual system components.
+Static power is measured when the system is inactive, while dynamic power is measured during the training of a ViT-B model.
+During training, power usage surges up to 6.5$\times$, with the GPU contributing 60% to the dynamic power.
+H.2 REP vs Miro
+In this experiment, we explore energy-accuracy trade-offs of REP-L and Miro.
+Figure 8 visually represents the comparison results.
+In each graph, the x-axis signifies total energy usage (lower is better), while the y-axis signifies final average accuracy (higher is better).
+Hence, a more cost-effective strategy is positioned closer to the upper-left corner of the graph.
+Regarding Miro, we maintain the hyperparameters as suggested in the original work [23] but incorporate the use of pre-trained ResNet-18 (RN18; 11M), ResNet-34 (RN34; 22M), and ResNet-50 (RN50; 25M)[11] models instead of non-pre-trained ones to enhance overall performance.
+For REP-L, we vary the number of training iterations per task insertion to match the energy usage of Miro variants with different ResNet models.
+Table 12: Comparison of REP-L with ADAM variations.
+Method Acc. ($\uparrow$) GPU Time (s) Mem. (GB)
+REP-L (ours) 72.58 2542.43 4.45
+VPT-Deep-L 66.70 1444.62 12.61
+VPT-Shallow-L 64.53 1271.69 6.89
+SSF-L 72.16 1573.70 22.09
+Adapter-L 62.05 680.05 12.01
+Table 13: Results on Split ImageNet-R, with ViT-L as the backbone model.
 Method Acc. ($\uparrow$) Fgt. ($\downarrow$) GPU Time (s) Mem. (GB)
-REP (ours) 75.34$\pm$0.16 3.58$\pm$0.29 2542.43 4.45
-L2P 61.57$\pm$0.66 9.73$\pm$0.47 1609.88 14.20
-DualPrompt 68.13$\pm$0.49 4.68$\pm$0.20 1507.42 12.30
-CODA-Prompt 75.45$\pm$0.56 1.64$\pm$0.10 1853.54 18.83
-Table 5: Results on Split ImageNet-R, with ViT-B as the backbone model (following the original setups).
-Competing Methods with Original Setups.
-For L2P, DualPrompt, and CODA-Prompt, we used the hyperparameters of the original papers, except for the number of iterations per task and batch size.
-In particular, smaller batch sizes must be adopted to fit into on-device memory constraints.
-Nonetheless, we present the results of three competing methods on Split ImageNet-R using their original hyperparameters in Table 5.
-Despite using the ViT-B backbone, they consume significantly more memory than REP due to their large batch sizes.
-5 Discussion
-Since our study focuses on scenarios wherein a sequence of tasks is presented within learning environments, our method does not determine the curriculum of learning.
-In general, actively selecting such curricula could enhance the final average accuracy, motivating future work to extend our method for curriculum-based continual learning.
-Moreover, although using replay buffer may not be feasible in real-world scenarios due to data privacy concerns, on-device CL is already viewed as a privacy-preserving approach, making rehearsal-based REP a promising direction.
+REP-L (ours) 75.34$\pm$0.16 3.58$\pm$0.29 2542.43 4.45
+CODA-Prompt-L 74.43$\pm$0.63 3.65$\pm$0.75 6134.51 15.93
+OVOR-L 75.40$\pm$0.74 4.17$\pm$0.65 5040.91 7.48
+REP-CODA 73.30$\pm$0.93 4.64$\pm$0.89 5459.71 13.71
+REP-OVOR 73.99$\pm$0.84 4.94$\pm$0.94 4486.41 7.48
+Figure 8: Energy-accuracy trade-offs for Split CIFAR100 and Split ImageNet-R between REP-L and Miro.
+To provide a spectrum of energy-accuracy trade-offs, we use pre-trained ResNet-18 (RN18), ResNet-34 (RN34), and ResNet-50 (RN50) for Miro.
+When operating within the same energy budget, REP-L consistently outperforms Miro variants, achieving 22–33% higher accuracy across datasets.
+REP-L demonstrates superior cost-effectiveness, especially on Split ImageNet-R compared to Split CIFAR-100.
+Both REP-L and Miro prove to be memory-efficient to some extent as they fit comfortably within the on-device memory capacity.
+Although we explore larger ResNet models for Miro, we do not observe the accuracy levels comparable to REP-L for either dataset.
+This experiment also underscores the importance of specifically optimizing vision transformers for on-device CL scenarios to advance performance boundaries.
+Appendix I Impact of Varying Hyperparameter Values
+Guided by the hyperparameter tuning strategies used in DualPrompt [38] and CODA-Prompt [31], we fine-tune key hyperparameters for our ALD approach using REP-L.
+Our focus is primarily on optimizing the threshold parameter ($\alpha$) and the adjustment factor ($\tau$), which are crucial for ALD’s adaptability and efficiency.
+We conduct a hyperparameter search using cross-validation, for which we designate 20% of our training dataset as a validation set.
+Figure 10 summarizes the results.
+We set the AToM’s merging parameter ($n$) to 8.
+We also explore the impact of varying prompt-related hyperparameters, as detailed in Figure 10.
+Note that computational costs increase with larger prompt sizes and longer prompt lengths, and higher costs do not always translate into improved accuracy.
+Based on Figure 10, we opt for a prompt pool size of 10 and a prompt length of 5 to strike the right balance between computational efficiency and model accuracy.
+Figure 9: Result for ALD’s hyperparameter tuning via cross-validation.
+Figure 10: Impact of prompt-related hyperparameters of REP. Note that when we increase the pool size and prompt length, there is a notable increase in computation cost.
+Appendix J Additional Related Works
+J.1 Adapter-based Rehearsal-free Methods
+Considering ADAM [44] as a state-of-the-art method for adapter-based CL, we evaluate all its variations proposed in [44] on Split ImageNet-R with 20 tasks and compare their final average accuracies with REP-L in Table 12.
+ADAM significantly reduces end-to-end training costs using an adapter model.
+Unlike REP-L, which performs the backward pass to update prompts, ADAM only updates the FC layers without the backward pass.
+However, training such a model is memory-intensive and often exceeds on-device memory capacity.
+Also, when implemented with the pre-trained ViT-L model, ADAM’s capability for generalization seems to be somewhat restricted.
+This observation does not necessarily indicate a fundamental limitation of ADAM but rather underscores the complexities of using it effectively in its current form.
+J.2 Advantages of REP over NAS
+While NAS (Neural Architecture Search) can be integrated into REP to drop some layers, it usually involves searching the model from scratch, which is time-consuming.
+In contrast, our method applies adaptive schedules directly to an existing pre-trained model, significantly saving computation time.
+So, it can naturally adapt to device-internal resource conditions without relying on external server resources.
+J.3 Online or Few-shot CL
+In on-device CL, the memory capacity for storing data from new tasks may be limited.
+Thus, one might consider adopting online or few-shot CL methods to save memory costs, as they require fewer new-task samples to be maintained in memory.
+Offline CL (our setting) typically yields better accuracy with higher resource usage.
+However, memory allocation for new samples is not a major constraint in offline CL.
+All prompt-based methods require 155–269MB to store new samples for both Split CIFAR-100 and Split ImageNet-R, accounting for only 3.4–5.9% of REP’s memory usage.
+Moreover, modern DL frameworks can store data in storage and proactively prefetch upcoming mini-batches for training, virtually eliminating memory concerns for new samples.
+J.4 Prompting in Self-Supervised CL
+Although our REP framework is designed for conventional supervised CL, we believe it can naturally extend to self-supervised prompt-based CL methods like HiDe-Prompt [34].
+By combining REP ’s resource-efficient techniques with hierarchical prompt optimization, this integration could enable on-device CL for large unlabeled datasets.
+Such a hybrid approach would preserve the performance benefits of prompt tuning while significantly reducing computational and memory overhead, making self-supervised CL more feasible on resource-constrained devices.
+Appendix K REP on Diverse Prompting Methods.
+In this experiment, we evaluate the performance and applicability of REP in approaches that directly update a single prompt without relying on prompt selection.
+CODA-Prompt [31] and OVOR [14] stand out as representative methods.
+We present experimental results using the Split CIFAR-100 and Split ImageNet-R datasets with a ViT-L backbone in  Table 13.
+REP offers moderately improved resource efficiency when applied to CODA-Prompt and OVOR.
+During the prompt update, we were able to apply ALD without mod-ifications, whereas AToM ’s main logic required reimplementation to function with CODA-Prompt and OVOR.
+This does not indicate a fundamental limitation of REP ’s core design but highlights the need to address this implementation challenge in future work.
+Nonetheless, we note that REP ’s baseline accuracy and resource efficiency (the first row in  Table 13) significantly outperform those observed in CODA-Prompt and OVOR.